OpenStudy (anonymous):

Question 2: Find the surface area of a rectangular prism with length of 10 in, width of 6 in, and height of 5 in.

OpenStudy (anonymous):

Surface area of rectangular prism=2ab + 2bc + 2ac. a, b, and c are the lengths of the 3 sides.

OpenStudy (anonymous):

A=2(wl+hl+hw) A=2(6*10+5*10+5*6) A=280